# Regge poles, grey body factors, and absorption cross sections for black hole metrics with discontinuity

### v1

@@ I Introduction
Characterizing black hole dynamics through gravitational wave observables has entered a critical phase of quantitative analysis with advanced interferometric detectors. A crucial topic to this endeavor is black hole spectroscopy, which involves modeling of ringdown spectral amplitudes, typically dominated by quasinormal modes (QNMs) [1, 2, 3, 4].
@@
As first pointed out by Leaver [5, 6], the QNMs can be attributed to the poles of the frequency-domain Green’s function of the underlying master equation that governs the dynamics of the perturbations, whereas the late-time tail is associated with the branch cut on the imaginary frequency axis [7]. The information inferred from the analytic properties of the Green’s function provides a unique avenue for understanding the characteristics of the resulting waveforms. Recently, the notion of spectral instability has aroused much attention. As pioneered by Nollert and Price [8, 9] and Aguirregabiria and Vishveshwara [10, 11], it was shown that even insignificant perturbations, such as a minor discontinuity [12, 13], can qualitatively deform the higher overtones in the QNM spectrum. This concept was elaborated and primarily established by Jaramillo et al. [14, 15] by systematically analyzing the impact on the QNM spectrum of randomized and sinusoidal perturbations to the metrics, and demonstrating the prone to instability, particularly against the ultraviolet, namely, high-frequency perturbations.
@@
These findings challenge the conventional assumption that an insignificant modification of the effective potential shall not introduce a sizable deviation in the resulting QNMs. Conversely, even in the presence of a rather moderate discontinuity [12, 13, 16], the asymptotic behavior of the QNM spectrum could be non-perturbatively modified. Specifically, high-overtone modes usually lie parallel along the real frequency axis rather than ascending the imaginary frequency axis observed for most black hole metrics [17, 18]. This phenomenon persists regardless of the discontinuity’s distance from the horizon or its magnitude. Specifically, the study of spectral instability and its ramifications has significant observational implications, especially for black hole spectroscopy [19, 20, 21, 22, 23, 24]. In real-world astrophysical contexts, gravitational radiation sources such as black holes or neutron stars are seldom isolated; they are typically submerged and interacting with surrounding matter. The resulting deviations from the ideal symmetric metrics cause the ringdown gravitational waves associated with the QNMs to differ substantially from those predicted for a pristine, isolated, compact object.
@@
The related topic of spectral instability, echoes, and causality has been explored extensively in recent years by many authors [25, 15, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Notably, Cheung et al. [27] pointed out that even the fundamental mode can be destabilized under generic perturbations. Specifically, a small perturbation to the Regge-Wheeler effective potential causes the fundamental mode to spiral outwardly while the deviation’s magnitude increases. Such an observation undermines the understanding that spectral instability might not significantly impact black hole spectroscopy as the fundamental mode is not subjected to spectral instability, leading to substantial observational implications. The physical significance of such instability has been further scrutinized, inclusively regarding a different scale introduced to the problem via the metric perturbations, in more recent studies [40, 41, 42, 43, 44].
@@
The present study is motivated by the above considerations. To this end, we generalize the matrix method to evaluate the Regge poles in black hole metrics with discontinuities. The obtained Regge pole spectrum is used to calculate the scattering amplitude and cross-section. Three key results emerge:
@@ II Generalized matrix method for Regge poles
where the spatial coordinate is known as the tortoise coordinate, and the effective potential is governed by the given spacetime metric, spin , and angular momentum with multipole number of the waveform. For instance, the Regge-Wheeler potential for the Schwarzschild black hole metric is
where
@@
The Regge poles [48, 49] are defined as the poles of the reflection amplitudes in the analytically continued angular momentum space, evaluated at a given (real-valued) frequency. Its resemblance to the QNMs is readily recognized as the above condition implies the boundary condition Eq. (8) as the reflection wave overwhelms the incident one. Subsequently, it leads to a vanishing Wronskian. Because the asymptotic waveforms given by Eqs. (5) and (6) are linearly dependent, a pole in thus implies a pole in and vice versa111It is worth noting that the flux conservation condition is no longer valid for complex angular momentum, which also invalidates Eq. (7).. Similar to the QNMs, it is usually assumed in the literature that these poles are simple ones, and one denotes the residue for the th pole as
@@
The differential cross-section is governed by the scattering amplitude
@@
with
The background contributions and will be specified below.
@@
The elements of the coefficient matrix are polynomials in of finite order. Specifically, this allows us to reiterate the matrix equation in the form
where the highest power in all the polynomials is for the Regge-Wheeler effective potential Eq. (2). Although Eq. (28) can be solved using the standard numerical methods, it is more convenient to transform it into a linear form [77] in . This procedure leads to a matrix equation for a generalized eigenvalue problem, which is readily solved using Mathematica’s subroutines Eigenvalue or Eigensystem.
@@ III Regge poles in modified Regge-Wheeler potentials
Eq. (36) furnishes a simplified scenario that is physically pertinent, corresponding to a mass shell wrapping around the black hole at .
@@ V Further discussions and concluding remarks
Combining these findings leads to a better understanding of why greybody factors are more stable observables in terms of Regge poles. On the one hand, the present study indicates that these modes become unstable only when the frequency is significant. On the other hand, the validity of the eikonal limit implies that the corrections from the Regge poles, particularly the high-order ones, become minor at higher frequencies. As a result, the overall effect from those unstable modes is suppressed, leading to more stable observables.

### v3

@@ I Introduction
Characterizing black hole dynamics through gravitational wave observables has entered a critical phase of quantitative analysis with the advent of interferometric detectors. A crucial topic to this endeavor is black hole spectroscopy, which involves modeling of ringdown spectral amplitudes, typically dominated by quasinormal modes (QNMs) [1, 2, 3, 4, 5].
@@
As first pointed out by Leaver [6, 7], the QNMs can be attributed to the poles of the frequency-domain Green’s function of the underlying master equation that governs the dynamics of the perturbations, whereas the late-time tail is associated with the branch cut on the imaginary frequency axis [8]. The information inferred from the analytic properties of the Green’s function provides a unique avenue for understanding the characteristics of the resulting waveforms. Recently, the notion of spectral instability has aroused much attention. As pioneered by Nollert and Price [9, 10] and Aguirregabiria and Vishveshwara [11, 12], it was shown that even insignificant perturbations, such as a minor discontinuity [13, 14], can qualitatively deform the higher overtones in the QNM spectrum. This concept was elaborated and primarily established by Jaramillo et al. [15, 16] by systematically analyzing the impact on the QNM spectrum of randomized and sinusoidal perturbations to the metrics, and demonstrating the prone to instability, particularly against the ultraviolet, namely, small-scale or high-frequency perturbations.
@@
These findings challenge the conventional assumption that an insignificant modification of the effective potential shall not introduce a sizable deviation in the resulting QNMs. Conversely, even in the presence of a rather moderate discontinuity [13, 14, 17], the asymptotic behavior of the QNM spectrum could be non-perturbatively modified. Specifically, high-overtone modes usually lie parallel along the real frequency axis rather than ascending the imaginary frequency axis observed for most black hole metrics [18, 19]. This phenomenon persists regardless of the discontinuity’s distance from the horizon or its magnitude. Specifically, the study of spectral instability and its ramifications has significant observational implications, especially in the context of black hole spectroscopy [20, 21, 22, 23, 24, 25]. In real-world astrophysical contexts, gravitational radiation sources such as black holes or neutron stars are seldom isolated; they are typically submerged and interacting with surrounding matter. The resulting deviations from the ideal symmetric metrics cause the ringdown gravitational waves associated with the QNMs to differ substantially from those predicted for a pristine, isolated, compact object.
@@
The related topic of spectral instability, echoes, and causality has been explored extensively in recent years by many authors [26, 16, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. Notably, Cheung et al. [28] pointed out that even the fundamental mode can be destabilized under generic perturbations. Specifically, a small perturbation to the Regge-Wheeler effective potential causes the fundamental mode to spiral outwardly while the deviation’s magnitude increases. Such an observation undermines the understanding that spectral instability might not significantly impact black hole spectroscopy. This is because it was understood that the fundamental mode is not subject to spectral instability. The fact that the fundamental mode might not be stable can have substantial observational implications [16, 29, 41]. The physical significance of such instability has been further scrutinized, inclusively regarding an interplay between the distance of the perturbation from the black hole and the size of the physical system, in more recent studies [42, 43, 44, 45, 46].
@@
The present study is motivated by the above considerations. To this end, we generalize the matrix method [57, 58] to evaluate the Regge poles in black hole metrics with discontinuities. The obtained Regge pole spectrum is used to calculate the scattering amplitude and cross-section. Three key results emerge:
@@ II Generalized matrix method for Regge poles
where the spatial coordinate is known as the tortoise coordinate, and the effective potential is governed by the given spacetime metric. For instance, the Regge-Wheeler potential for the Schwarzschild black hole metric is [59]
where is the spin and is the multipole number of the angular momentum of the perturbations. Also,
@@
The Regge poles [50, 51] are defined as the poles of the reflection amplitudes in the analytically continued angular momentum space, evaluated at a given (real-valued) frequency while (formally) adopting the same boundary condition Eq. (8). Its resemblance to the QNMs is readily recognized as the boundary condition implies divergence in the scattering matrix as the reflection wave overwhelms the incident one. Subsequently, it leads to a vanishing Wronskian. Because the asymptotic waveforms given by Eqs. (5) and (6) are linearly dependent, a pole in thus implies a pole in and vice versa111It is worth noting that the flux conservation condition is no longer valid for complex angular momentum, which also invalidates Eq. (7).. Similar to the QNMs, it is usually assumed in the literature that these poles are simple ones, and one denotes the residue for the th pole as
@@
with
Here, and are evaluated at th Regge pole. The order of the Legendre polynomials , inherited from the partial wave formula, is also governed by the Regge poles as a result of the contour integration. The differential cross-section is given by the scattering amplitude
@@
The background contributions and come from the high-energy limit of the scattering process and can be estimated using means such as the WKB approximation (also see further discussions in Sec. IV).
@@
We note that Eq. (28) is precisely the same equation proposed in the original matrix method [57], except that we are now solving for unknown complex angular momenta in place of frequency. This was facilitated by the fact that the form of the boundary condition Eq. (8) remains unchanged. The elements of the coefficient matrix are polynomials in of finite order. Specifically, this allows us to rewrite the matrix equation in the form
where all polynomials are at most quadratic in for the Regge-Wheeler effective potential Eq. (2). Although Eq. (28) can be solved using the standard numerical methods, it is more convenient to transform it into a linear form [80] in . This procedure leads to a matrix equation for a generalized eigenvalue problem, which is readily solved using Mathematica’s subroutines Eigenvalue or Eigensystem.
@@ III Regge poles in modified Regge-Wheeler potentials
Here, for simplicity, we have started from the effective potential, and the corresponding metrics associated with Eqs. (34)-(36) should be derived by solving for the unknown function in some metric ansatz. However, due to the simple forms, one can make use of Birkhoff’s theorem for spherically symmetric cases and conclude that Eq. (36) furnishes a simplified scenario that is physically pertinent, corresponding to an infinitely thin mass shell wrapping around the black hole. For the present case, we understand that when the thickness of the mass shell becomes finite, it does not lead to a significant modification of the effective potential that possibly triggers the spectral instability in QNMs. Nonetheless, the physical significance of perturbations to the metric and effective potential is a crucial topic; a small modification in the effective potential does not necessarily imply that the corresponding metric perturbation is insignificant. These concerns have been elaborated recently in [42].
@@ V Further discussions and concluding remarks
Combining these findings leads to a better understanding of why greybody factors are more stable observables in terms of Regge poles. On the one hand, the present study indicates that these modes become unstable only when the frequency is significant (for the setup employed in the present study, this occurs at ). On the other hand, the validity of the eikonal limit implies that the corrections from the Regge poles, particularly the high-order ones, become minor at higher frequencies. As a result, the overall effect from those unstable modes is suppressed, leading to more stable observables.